OpenStudy (anonymous):

Two particles having charges of 0.570nC and 20.5nC are separated by a distance of 2.00m. A) At what point along the line connecting the two charges is the net electric field due to the two charges equal to zero? B) Where would the net electric field be zero if one of the charges were negative?

OpenStudy (whpalmer4):

electrostatic force varies inversely with the square of the distance and directly with the charge. Distance from one end to the point of interest is r, and the distance from the other end is 2.00-r. \[\frac{0.570}{r^2} = \frac{20.5}{(2.00-r)^2}\]Cross-multiply and solve for \(r\).